Suppose that a 2.00L glass flask is 0.500 mol of H2 and 0.500 mol Cl2 are reacted to produce HCl. When equilibrium is reached, the concentrations of HCl is found to be 0.110M. What is the keq? (ICE box, Keq expression & Keq value)
Answer #1

H2 + Cl2 \leftrightharpoons 2HCl

ICE box
H2 Cl2 2HCl
Initial 0.5/2 0.5/2 0
Change -x -x 2x
Equilibrium 0.5/2-x 0.5/2-x 0.11

Keq = [HCl]2 /[H2][Cl2]

[HCl] = 0.11 = 2x therefore ,x = 0.11/2 = 0.055

[H2] = 0.5/2 - 0.055 = 0.195

[Cl2] = 0.5/2 - 0.055 = 0.195

Keq = (0.055)2/(0.195)2 = 0.079